# Sequential model confidence sets

### v1

@@ Abstract
In most prediction and estimation situations, scientists consider various statistical models for the same problem, and naturally want to select amongst the best. Hansen et al. (2011) provide a powerful solution to this problem by the so-called model confidence set, a subset of the original set of available models that contains the best models with a given level of confidence. Importantly, model confidence sets respect the underlying selection uncertainty by being flexible in size. However, they presuppose a fixed sample size which stands in contrast to the fact that model selection and forecast evaluation are inherently sequential tasks where we successively collect new data and where the decision to continue or conclude a study may depend on the previous outcomes. In this article, we extend model confidence sets sequentially over time by relying on sequential testing methods. Recently, e-processes and confidence sequences have been introduced as new, safe methods for assessing statistical evidence. Sequential model confidence sets allow to continuously monitor the models’ performances and come with time-uniform, nonasymptotic coverage guarantees.
Keywords: Model confidence set, forecast evaluation, model selection, sequential inference, multiple testing.
@@ 1 Introduction
In science as well as in our daily life, we frequently encounter situations in which multiple statistical models or forecasts are available for the same problem and where we have to decide which model(s) or forecast(s) we want to trust. We might think for example of multiple inflation regression models with respect to different sets of covariates or of multiple weather institutions which issue precipitation predictions for the next day. In such situations, we naturally want to select the best model(s) or forecast(s), where the term “best” is defined in terms of a user-specified criterion.
@@
The model confidence set (MCS) proposed by Hansen et al. (2011) provides a promising solution to this problem. It departs from standard practice where just a single model is selected according to some appropriate loss and nothing is said about the uncertainty associated with this selection. The MCS takes this uncertainty into account by reducing the original set of available models to a smaller set of flexible size that contains the best models with a given level of confidence. Model confidence sets are of great importance in applications without an obvious benchmark, may be easily constructed by the MCS algorithm, and are widely applied in the econometrics community; see, e.g., Weron (2014), Corbet et al. (2019) or Masini et al. (2023) amongst others. However, even though the MCS provides an appealing solution to a highly important problem, it comes with limitations. In particular, it is assumed that the sequences of observed loss differences are stationary, an assumption which may be questionable in practise since we may expect that the models evolve over time, and correct for errors and systematic biases by using past information. Apart from relying on strong stationarity assumptions, the MCS requires a sample of some fixed size chosen independently of the data. In other words, we assume that we observe losses of the models over some prespecified evaluation period in order to compute the MCS only once at the end of the period. This procedure highly contrasts our natural urge to assess the models on a regular basis by successively including past observations to the available dataset. Consider for example different weather institutions that predict the accumulated precipitation for the following day on a daily-basis. Then, firstly, it is not given that the initially better models remain superior until the end of our study. Secondly, we want to assess the forecasters and monitor their performance sequentially, say, at the end of each day, week or month, since the forecasters issue their forecasts sequentially as well and may behave nonstationary. If we gather evidence that one institution is worse than the others but the evidence is not yet statistically significant, then we want to continue to collect evidence for this hypothesis without loosing the information of the earlier observations. Crucially, we want to decide at the end of each day, week or month whether to continue or conclude the study, depending on the previous outcomes.
@@
In summary, model selection and forecast evaluation are inherently sequential tasks and we should seek for methods that allow to sequentially choose amongst the best models while respecting the underlying selection uncertainty. In this article, we address this need and contribute with sequential model confidence sets. Sequential model confidence sets generalize the core idea of Hansen et al. (2011) sequentially over time and enjoy time-uniform coverage guarantees of the theoretically superior objects. Our methods rely on e-processes and time-uniform confidence sequences, which lay the foundation for safe anytime-valid inference, a field which met a surge of new contributions over the last years by various authors; see, e.g., Shafer (2021), Vovk and Wang (2021), Ramdas et al. (2023) and Grünwald et al. (2019) amongst others.
@@
Sequential testing methods rely on minimal distributional assumptions and have been used before for sequential forecast evaluation: Henzi and Ziegel (2022) provide e-processes to sequentially assess two probability forecasts for the strong hypothesis that one forecaster is better than the other at all time points. In contrast, Choe and Ramdas (2023) contribute with e-processes for the weaker hypothesis that one forecaster is uniformly better than the other on average. Both studies assess the performance of some forecaster at time $t\in\mathbb{N}$ with respect to the (average) expected loss difference given all information until the previous time step, which goes back to the test of equal conditional predictive ability proposed by Giacomini and White (2006). In contrast to the unconditional approach, most prominently in the seminal work of Diebold and Mariano (1995), the conditional approach uses available information to infer which forecaster is more accurate at a specific date. In this article, we follow the conditional approach and study the strong as well as the weak hypothesis which we further divide into two subhypotheses. In contrast to all aforementioned contributions on forecast comparison, our methods allow to simultaneously draw inference on the performance of $m\geq 2$ different forecasters rather than just being able to compare two different forecasters.
@@
Although the article mainly focuses on forecast evaluation, many of the presented results directly extend to model selection as long as the used information criterion satisfies the imposed assumptions. For this reason, we introduce the problem in a general framework and illustrate the connections to forecast comparison and model selection afterwards. In order to construct sequential model confidence sets, we make use of the e-processes provided by Choe and Ramdas (2023) under the assumption that the observed loss differences are conditionally bounded, which does not only hold for bounded losses but also for prominent choices of scoring rules such as the continuous ranked probability score (CRPS; Matheson and Winkler, 1976) or the quantile score (Gneiting, 2011).
@@
The remainder of the paper is structured as follows. Section 2 presents the problem and introduces sequential model confidence sets. In Section 3, we review important definitions from sequential testing theory and provide methods to construct sequential model confidence sets with respect to three particular notions of superior objects. A simulation study is conducted in Section 4 before we apply the proposed methods in case studies on Covid-19 related deaths and wind gust predictions in Section 5. The main part of the paper closes with a discussion in Section 6. A review on forecast evaluation, technical comments, and additional information for the case studies are available in Appendices A through G.
@@ 2 Sequential model confidence sets
We consider a set $\mathcal{M}_{0}=\{1,\dots,m\}$ indexing $m\geq 2$ different statistical models or forecasters, simply referred to as models, and let $L$ be a loss function which measures the quality of the models. We let $L$ be negatively oriented, that is, lower scores correspond to better performance, assume that we collect data in discrete time and denote the loss of model $i\in\mathcal{M}_{0}$ at time $t\in\mathbb{N}$ by $L_{i,t}$. We assume that all random quantities are defined on some underlying measurable space $(\Omega,\mathcal{F})$ equipped with some filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. We denote the family of all probability measures on $(\Omega,\mathcal{F})$ by $\mathfrak{B}(\Omega)$ and write $\mathbb{Q}$ for a generic element in $\mathfrak{B}(\Omega)$.
Example 2.1 (Sequential forecasters).
@@
Let $\{1,\dots,m\}$ denote different forecasters which all sequentially issue predictive distributions or point forecasts $(f_{1,t})_{t\in\mathbb{N}},\dots,(f_{m,t})_{t\in\mathbb{N}}$ for some unknown quantity $(Y_{t})_{t\in\mathbb{N}}\subseteq\mathbb{R}$. For each $i=1,\dots,m$ and $t\in\mathbb{N}$, the forecast $f_{i,t}$ is predictable and refers to the observation $Y_{t}$. To assess the forecasters, one should employ proper scoring rules or consistent scoring functions, see Appendix A. Then, the time series $(L_{i,t})_{t\in\mathbb{N}}$ emerges by applying a proper scoring rule (consistent scoring function) $L$ to the forecast-observation pairs $(f_{i,t},Y_{t})_{t\in\mathbb{N}}$, that is $L_{i,t}=L(f_{i,t},Y_{t})$, $i=1,\dots,m,t\in\mathbb{N}$.
Example 2.2 (Sequential model selection).
Let $\{1,\dots,m\}$ consist of nested and non-nested regression models for some sequentially revealed response $(Y_{t})_{t\in\mathbb{N}}$. At each $t\in\mathbb{N}$, model $i$ observes the covariate(s) $\boldsymbol{X}_{i,t}\subseteq\mathbb{R}^{p_{i}}$, $p_{i}\geq 1,$ and proclaims $Y_{t}=\boldsymbol{\beta}_{i,t}^{\top}\boldsymbol{X}_{i,t}+\varepsilon_{i,t}$ for some $\boldsymbol{\beta}_{i,t}\subseteq\mathbb{R}^{p_{i}}$ and random error $\varepsilon_{i,t}$. The quality of the models is typically evaluated with respect to some likelihood-based information criterion (Ding et al., 2018). For example, for ^$\hat{\ell}_{i,t}$ the empirical log-likelihood of model $i$ given all observations until $t$, $L_{i,t}=-2\hat{\ell}_{i,t}+2p_{i}$ or $L_{i,t}=-2\hat{\ell}_{i,t}+p_{i}\log t$ reveal the well-known Akaike information criterion (AIC) or Bayesian information criterion (BIC), respectively.
Next, we clarify the notion of “best models at time $t\in\mathbb{N}$. For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, consider
$d_{ij,t}=L_{i,t}-L_{j,t},\quad\mu_{ij,t}=\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})$
and their averages
$\hat{\Delta}_{ij,t}=\frac{1}{t}\sum_{s=1}^{t}d_{ij,s},\quad\Delta_{ij,t}=\frac{1}{t}\sum_{s=1}^{t}\mu_{ij,s},$
@@
where we omit the dependence of $\mu_{ij,t}$ and $\Delta_{ij,t}$ on the choice of the probability measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$ for the sake of clarity. Whereas Diebold and Mariano (1995) define model (forecast) superiority with respect to the unconditional expected loss differences, we follow Giacomini and White (2006), Lai et al. (2011), Henzi and Ziegel (2022) and Choe and Ramdas (2023) by defining the superior models in terms of the (average) conditional expected loss differences. In particular, we consider the strongly superior objects
$\mathcal{M}^{\textrm{s},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij,r}\leq 0\textrm{ for all }r\leq t,\textrm{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N},$ (1)
@@
$\mathcal{M}^{\textrm{uw},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,r}\leq 0\textrm{ for all }j\in\mathcal{M}_{0},\textrm{ for all }r\leq t\right\},\quad t\in\mathbb{N},$ (2)
@@
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,t}\leq 0\textrm{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$ (3)
For each $t\in\mathbb{N}$, $\mathcal{M}^{\textrm{s},\star}_{t},\mathcal{M}^{\textrm{uw},\star}_{t}$ and $\mathcal{M}^{\textrm{w},\star}_{t}$ are $\mathcal{F}_{t-1}$-measurable random sets with $\mathcal{M}^{\textrm{s},\star}_{t}\subseteq\mathcal{M}^{\textrm{uw},\star}_{t}\subseteq\mathcal{M}^{\textrm{w},\star}_{t}$. Importantly, $\mathcal{M}^{\textrm{w},\star}_{t}\neq\emptyset$ for all $t\in\mathbb{N}$, whereas $\mathcal{M}^{\textrm{s},\star}_{t}$ and $\mathcal{M}^{\textrm{uw},\star}_{t}$ may eventually become empty for all $t\geq t_{0}$ for some $t_{0}\in\mathbb{N}$.
Let $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ be the targeted sequence of superior objects. Of particular interest for this article are the sequences given at (1), (2) and (3), but one could also think of other possible notions of best models. For a given confidence level $\alpha\in(0,1)$, we call (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a sequence of model confidence sets or sequential model confidence sets, for short SMCS(s), for $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}$ at level $\alpha$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}\left(\exists t\geq 1:\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t}\right)\leq\alpha,$ (4)
or equivalently, if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$,
$\mathbb{Q}\left(\forall t\geq 1:\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t}\right)\geq 1-\alpha.$ (5)
We refer to the equivalent conditions (4) and (5) as time-uniform coverage guarantees of the sequential model confidence sets and highlight the fact that the time quantifiers are inside of the probability, which is a much stronger requirement than assuming $\mathbb{Q}(\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t})\geq 1-\alpha$ for all $t\in\mathbb{N}$ (or assuming that there exists a $t\in\mathbb{N}$ with $\mathbb{Q}(\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$); see, e.g., Howard et al. (2021, Lemma 3).

### v3

@@ Abstract
In most prediction and estimation situations, scientists consider various statistical models for the same problem, and naturally want to select amongst the best. Hansen et al. (2011) provide a powerful solution to this problem by the so-called model confidence set, a subset of the original set of available models that contains the best models with a given level of confidence. Importantly, model confidence sets respect the underlying selection uncertainty by being flexible in size. However, they presuppose a fixed sample size which stands in contrast to the fact that model selection and forecast evaluation are inherently sequential tasks where we successively collect new data and where the decision to continue or conclude a study may depend on the previous outcomes. In this article, we extend model confidence sets sequentially over time by relying on sequential testing methods through e-processes and confidence sequences. Sequential model confidence sets allow to continuously monitor the models’ performances and come with time-uniform, nonasymptotic coverage guarantees.
Keywords: Model confidence set, forecast evaluation, forecast comparison, sequential inference, multiple testing.
@@ 1 Introduction
In science as well as in our daily life, we frequently encounter situations in which multiple statistical models or forecasts are available for the same problem and where we have to decide which model(s) or forecast(s) we want to trust. We might think for example of multiple regression models for inflation with respect to different sets of covariates or of multiple weather services which issue precipitation predictions for the next day. In such situations, we naturally want to select the best model(s) or forecast(s), where the term “best” is defined in terms of a user-specified criterion, typically given by some real-valued loss function where lower scores correspond to better performance.
The model confidence set (MCS) proposed by Hansen et al. (2011) provides a promising solution to this problem. It departs from standard practice where just a single model is selected according to some appropriate loss and nothing is said about the uncertainty associated with this selection. The MCS takes this uncertainty into account by reducing the original set of available models to a smaller set of flexible size that contains the best models with a given level of confidence. Model confidence sets are of great importance in applications without an obvious benchmark, may be easily constructed by the MCS algorithm, and are widely applied in the econometrics community; see, e.g., Weron (2014), Corbet et al. (2019) or Masini et al. (2023) amongst others.
@@
However, even though the MCS provides an appealing solution to a highly important problem, it comes with limitations. In particular, it is assumed that the sequences of observed loss differences are stationary, an assumption which may be questionable in practice since we may expect that the models evolve over time, and correct for errors and systematic biases by using past information. Apart from relying on strong stationarity assumptions, the MCS requires a sample of some fixed size chosen independently of the data. In other words, we assume that we observe losses of the models over some prespecified evaluation period in order to compute the MCS only once at the end of the period. This procedure highly contrasts our natural urge to assess the models on a regular basis by successively including past observations to the available dataset. Consider for example different weather institutions that predict the accumulated precipitation for the following day on a daily-basis. Then, firstly, it is not given that the initially better models remain superior until the end of our study. Secondly, we want to assess the forecasters and monitor their performance sequentially, say, at the end of each day, week or month, since the forecasters issue their forecasts sequentially as well and may behave nonstationarily. If we gather evidence that one institution is worse than the others but the evidence is not yet statistically significant, then we want to continue to collect evidence for this hypothesis without losing the information of the earlier observations. Crucially, we want to decide at the end of each day, week or month whether to continue or conclude the study, depending on the previous outcomes. Finally, the derivation of the MCS is based on bootstrap approximations and asymptotic, CLT-related properties. In addition to high computational costs, coverage guarantees are only asymptotic.
@@
Selecting models and comparing forecasts are inherently sequential tasks, and we should seek methods that allow for a sequential choice of the best models while respecting the underlying selection uncertainty. In this article, we address this need and contribute with sequential model confidence sets. Sequential model confidence sets generalize the core idea of Hansen et al. (2011) sequentially over time and enjoy time-uniform coverage guarantees of the theoretically superior objects. Moreover, they have non-asymptotic validity and do not rely on computationally expensive approximations. Our methods rely on e-processes and time-uniform confidence sequences, which lay the foundation for safe anytime-valid inference, a field which met a surge of new contributions over the last years by various authors; see, e.g., Shafer (2021), Vovk and Wang (2021), Ramdas et al. (2023) and Grünwald et al. (2024) amongst others.
@@
Sequential testing methods relying on minimal distributional assumptions have been used before for sequential forecast evaluation: Henzi and Ziegel (2022) provide e-processes to sequentially test the strong hypothesis that one forecaster is better than another at all time points. In contrast, Choe and Ramdas (2023) develop e-processes for the weaker hypothesis that one forecaster is uniformly better than the other on average. Both studies assess the performance of some forecaster at time $t\in\mathbb{N}$ with respect to the (average) expected loss difference given all information up to the previous time step. This is relevant to tests of equal conditional predictive ability proposed by Giacomini and White (2006). In contrast to the unconditional approach, most prominently in the seminal work of Diebold and Mariano (1995), the conditional approach uses the available information to infer which forecaster is more accurate on a specific date. In this article, we follow the conditional approach and study the strong as well as the weak hypothesis. In contrast to all aforementioned contributions on forecast comparison, our methods allow to simultaneously draw inference on the performance of $m\geq 2$ different forecasters rather than only being able to compare two different ones.
@@
In order to construct sequential model confidence sets, we make use of the e-processes provided by Howard et al. (2020) and Choe and Ramdas (2023) under the assumption of sub-Gaussian or sub-exponential loss differences, which subsumes in particular the case where the loss differences are conditionally bounded, an assumption, which does not only hold for bounded losses but also for prominent choices of scoring rules such as the continuous ranked probability score (CRPS; Matheson and Winkler, 1976) or the quantile score (Gneiting, 2011).
@@
The remainder of the paper is structured as follows. Section 2 presents the problem and introduces sequential model confidence sets. In Section 3, we review important definitions from sequential testing theory and provide methods to construct sequential model confidence sets with respect to three particular notions of superior objects. A simulation study is conducted in Section 4 before we apply the proposed methods in case studies on Covid-19 related deaths and wind gust predictions in Section 5. The main part of the paper closes with a discussion in Section 6. Some background on forecast evaluation, technical comments, and proofs are available in the supplementary material.
@@ 2 Sequential model confidence sets
We consider a set $\mathcal{M}_{0}=\{1,\dots,m\}$ indexing $m\geq 2$ different statistical models or forecasters, simply referred to as models, and we let $L$ be a loss function which measures the quality of the models and is negatively oriented, that is, lower values correspond to better performance. We collect data in discrete time and denote the loss of model $i\in\mathcal{M}_{0}$ at time $t\in\mathbb{N}$ by $L_{i,t}$. We assume that all random quantities are defined on some underlying measurable space $(\Omega,\mathcal{F})$ equipped with some filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. We denote the family of all probability measures on $(\Omega,\mathcal{F})$ by $\mathfrak{B}(\Omega)$ and write $\mathbb{Q}$ for a generic element in $\mathfrak{B}(\Omega)$.
@@
Our leading example of the above setting is the following. Let $1,\dots,m$ index different forecasters which all sequentially issue predictive distributions or point forecasts $(f_{1,t})_{t\in\mathbb{N}},\dots,(f_{m,t})_{t\in\mathbb{N}}$ for some unknown quantity $(Y_{t})_{t\in\mathbb{N}}$, adapted to $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. For each $i=1,\dots,m$ and $t\in\mathbb{N}$, the forecast $f_{i,t}$, which refers to the outcome $Y_{t}$, is predictable, meaning that it is only based on information until time $t-1$. To assess the forecasters, one should employ proper scoring rules or consistent scoring functions, see Supplement A for further details. Then, the time series $(L_{i,t})_{t\in\mathbb{N}}$ emerges by applying a proper scoring rule (consistent scoring function) $L$ to the forecast-observation pairs $(f_{i,t},Y_{t})_{t\in\mathbb{N}}$, that is $L_{i,t}=L(f_{i,t},Y_{t})$, $i=1,\dots,m$, $t\in\mathbb{N}$.
Next, we clarify the notion of “best models”. For $i,j\in\mathcal{M}_{0}$ and $t\in\mathbb{N}$, consider $d_{ij,t}=L_{i,t}-L_{j,t}$ and $\mu_{ij,t}=\mathbb{E}(d_{ij,t}\mid\mathcal{F}_{t-1})$ as well as their averages $\hat{\Delta}_{ij,t}=\sum_{s=1}^{t}d_{ij,s}/t$ and $\Delta_{ij,t}=\sum_{s=1}^{t}\mu_{ij,s}/t$, where we omit the dependence of $\mu_{ij,t}$ and $\Delta_{ij,t}$ on the choice of the probability measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$ for the sake of brevity.
@@
We follow Giacomini and White (2006), Lai et al. (2011), Henzi and Ziegel (2022) and Choe and Ramdas (2023) by defining the superior models in terms of the (average) conditional expected loss differences. In particular, we define the strongly superior objects
$\mathcal{M}^{\textrm{s},\star}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij,t}\leq 0\textrm{ a.s.~{}for all }j\in\mathcal{M}_{0},\,t\in\mathbb{N}\right\},$ (1)
@@
$\mathcal{M}^{\textrm{uw},\star}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,t}\leq 0\textrm{ a.s.~{}for all }j\in\mathcal{M}_{0},\,t\in\mathbb{N}\right\},$ (2)
@@
$\mathcal{M}^{\textrm{w},\star}_{t}=\left\{i\in\mathcal{M}_{0}\mid\Delta_{ij,t}\leq 0\textrm{ a.s.~{}for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$ (3)
For each $t\in\mathbb{N}$, we have $\mathcal{M}^{\textrm{s},\star}\subseteq\mathcal{M}^{\textrm{uw},\star}\subseteq\mathcal{M}^{\textrm{w},\star}_{t}$. Importantly, $\mathcal{M}^{\textrm{w},\star}_{t}\neq\emptyset$ for all $t\in\mathbb{N}$, whereas $\mathcal{M}^{\textrm{s},\star}$ and $\mathcal{M}^{\textrm{uw},\star}$ may be empty.
Let $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ be the targeted sequence of superior objects. In the case of (1) or (2), we just consider a constant sequence. For a given confidence level $\alpha\in(0,1)$, we call (^$(\widehat{\mathcal{M}}_{t})_{t\in\mathbb{N}}\subseteq\mathcal{M}_{0}$ a sequence of model confidence sets or sequential model confidence sets, for short SMCS(s), for $(\mathcal{M}^{\star}_{t})_{t\in\mathbb{N}}$ at level $\alpha$ if, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, $\mathbb{Q}(\exists t\geq 1:\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$, or equivalently
$\mathbb{Q}(\forall t\geq 1:\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t})\geq 1-\alpha.$ (4)
We refer to this property as a time-uniform coverage guarantee for the sequential model confidence sets and highlight the fact that the time quantifiers are inside the probability, which is a much stronger requirement than assuming $\mathbb{Q}(\mathcal{M}^{\star}_{t}\subseteq\widehat{\mathcal{M}}_{t})\geq 1-\alpha$ for all $t\in\mathbb{N}$ (or assuming that there exists a $t\in\mathbb{N}$ with $\mathbb{Q}(\mathcal{M}^{\star}_{t}\nsubseteq\widehat{\mathcal{M}}_{t})\leq\alpha$); see, e.g., Howard et al. (2021, Section 1.1), for a historical perspective on confidence sequences.
In the paper by Hansen et al. (2011), the authors make the assumption that the (unconditional) expectation $\mathbb{E}(d_{ij,t})=\mu_{ij}$ does not depend on $t$, and they define the set of superior objects as =$\mathcal{M}^{\star}=\left\{i\in\mathcal{M}_{0}\mid\mu_{ij}\leq 0\text{ for all}j\in\mathcal{M}_{0}\right\}.$ This is in line with Diebold and Mariano (1995), who also define model (forecast) superiority with respect to the unconditional expected loss differences. For a fixed sample size, they target an asymptotic guarantee of the form (^$\mathbb{Q}(\mathcal{M}^{\star}\subseteq\widehat{\mathcal{M}})\geq 1-\alpha$ for their model confidence set (MCS) $\widehat{\mathcal{M}}$. Our method provides a stronger guarantee but it is also the first method with finite sample validity for a fixed sample size.
Model confidence sets are also related to the literature on inference on the argmin of a vector of means, see, e.g. Futschik and Pflug (1995); Zhang et al. (2024). In this strand of literature, the desired coverage guarantee is typically weaker in that $\mathbb{Q}(i\in\widehat{\mathcal{M}})\geq 1-\alpha$ should hold for each superior model $i$ but not uniformly over all models.